OpenStudy (anonymous):

Find the length of the slant height of a cone with a radius of 15 cm and a surface area of 1884 square cm.

OpenStudy (anonymous):

surface area of cone= pi *r *l \[\pi \times 15 \times l=1884,l=\frac{ 1884 }{ 15 \pi }=\frac{ 1884 }{ 15 \times 3.14 }=\frac{ 1884\times 100 }{ 15 \times 314 } =40\] 

OpenStudy (anonymous):

?????

OpenStudy (mathstudent55):

\(\large A_{cone} = \pi r^2 + \pi r l\) Let's start with the surface of the base. \(\large A_{base} = \pi r^2 = \pi \times (15~cm)^2 = 707~ cm^2\) Subtract 707 cm^2 from the surface area, and you have the lateral surface area. Lateral surface area = 1884 cm^2 - 707 cm^2 = 1177 cm^2 The lateral surface area = \(\pi r l\), and r = 15 cm. \(\pi r l = 1177~cm^2\) \(3.14159 \times 15~cm \times l = 1177~cm^2\) Solve for \(l\)

OpenStudy (anonymous):

l=25

OpenStudy (mathstudent55):

Correct.

OpenStudy (anonymous):

yesssssssss